Spring 2019 Surveying 2302 SCORE SECT Name L- 425 +2.1% BVC 880.00 470.72 92 -3.6% EVC Low PointPLEASE HELPPPP
Answer #1

Answers:

It is an equal tangent vertical curve

L = 425 '

L/2 = 425/2 = 212.5 '

STA PVI  = STA BVC + (L/2) = 2880 + 212.5

= 3092.5 ft

= STA 30+92.5

Elevation PVI = Elev BVC - g1*(L/2)

= 470.72 - (3.6/100) * 212.5

= 463.07 ft

STA EVC = STA PVI + (L/2)

= 3092.5 + 212.5

= 3305 ft

= STA 33+05

Elev EVC = Elev PVI + g2*(L/2)

= 463.07 + (2.1/100) *212.5

= 467.5325

Low point

x = distance from BVC to low point

x = g1*L / (g1 - g2)

= (-3.6)* 425 / ( -3.6 - 2.1)

= 268.42'

LP STA = STA BVC + x

= 2880 + 268.42

= 3148.42

= STA 31 + 48.82

y = yBVC + g1*x + (r/2)*x^2

r /2= (g2 -g1)/ 2L

= ( 2.1 + 3.6) / 2*4.25

= 0.671

y LP = 470.72 - (3.6) *2.6842 + 0.671 *2.6842^2

= 465.89 ' = elevation of low point